A seconds pendulum is taken to planet where the acceleration due to gravity is 4 times that in earth .What would be the time period of the pendulum on the planet
 

Grade:9

1 Answers

Kruthic
6 years ago
Time period of a simple pendulum=2*pi*(l/g)1/2
where l is the length of the string 
and g is the gravition of the planet
Since it is a seconds pendulum , 2*pi*(l/g)1/2=1 second (in earth)
On the other planet , gravity = 4g
Therefore new time period= 2*pi*(l/4g)1/2 =2*pi*(l/g)1/2*(1/4)1/2
=1 second * ½ =0.5 seconds